OpenStudy (anonymous):

Can you see if this is right? For how many x values will the tanget lines to y=4sinx and y=x^2/2 be parallel?

OpenStudy (anonymous):

y'=4cosx 4cosx=1/2 cosx=1/8

OpenStudy (anonymous):

or is it 4cosx=x^2/2

OpenStudy (anonymous):

4cosx = x is the answer!

OpenStudy (anonymous):

for some reason the answer is supposed to be 3...

OpenStudy (kainui):

Ahh I see. Take the derivatives and graph them and see where they intersect. They intersect 3 points which means that the slopes of the graphs at that x-value is exactly the same.

OpenStudy (anonymous):

do i take the derivative of the second equation too?

OpenStudy (anonymous):

or just the first one?

OpenStudy (anonymous):

i just derived them and equal them. theie ramp should be parallel!

OpenStudy (anonymous):

so would it be 4cosx=x^2/2 or 4cosx=x?

OpenStudy (anonymous):

the second derive is X. \[^{2}=2x \rightarrow 2x.1/2=X\]

OpenStudy (anonymous):

can you explain that?

OpenStudy (anonymous):

clear?!

OpenStudy (anonymous):

whats your equation supposed to be?

OpenStudy (anonymous):

2x.1/2=x ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Kay thank you

OpenStudy (anonymous):

ur pleasure;) if u need more explains, i can help:)